OpenStudy (anonymous):

I need help solving: One number is 4 more than another. IF 6 times the smaller minus 4 times the larger is 4, what are the two numbers.

OpenStudy (anonymous):

We have simultaneous equations: \[ x = y + 4 \] and \[ 6y - 4x = 4 \]

OpenStudy (anonymous):

Thanks gentoolx. Ok I get that, are they solved separately, or is one solved before the other?

OpenStudy (anonymous):

So, you can just substitute the value from first equation (x value) to the second equation or you can transform the first to \[ x-y=4 \] and then: \[4x-4y=16\] Then you sum up these two equations to get \[2y = 20 \] Then you compute value of y and then you substitute it to one of the original equations to compute x.

OpenStudy (anonymous):

To be clear: when I meant "these two equations" I meant transformed equation and the second equation from original ones.

OpenStudy (anonymous):

Thanks again, I am sorry how do you get the 4 on the other side of the equal sign

OpenStudy (anonymous):

you multiplied 4 by both sides

OpenStudy (anonymous):

I multiplied both sides by four.

OpenStudy (anonymous):

\[x = y + 4\] \[x - y = 4 | * 4\]\[4x - 4y = 16\]

OpenStudy (anonymous):

More info about solving simultaneous equations here: http://en.wikipedia.org/wiki/Simultaneous_equations

OpenStudy (anonymous):

Thanks so much for your help, I will try that website, its been a long time since I took algebra

OpenStudy (anonymous):

You're welcome.